# Electromagnetic field in a cavity induced by gravitational waves

### v1

@@ V Summary and Discussion
We introduce physical laws in a globally hyperbolic spacetime described by Einstein’s equations. The contracted Bianchi identities imply the conservation of stress-energy. The congruence of the timelike geodesics without vorticity, have their spatial distribution determined by the extrinsic curvature. However, material elements, despite being influenced by gravity, deviate from geodesics due to elasticity stemming from their deformation. To account for the acoustic oscillation of the material, we introduce the 4-velocity and strain, both contributing the stress-energy. The equations of motion for these material elements encompass the temporal symmetry of the material metric, energy conservation, and the conservation of stress-energy. The EM field is governed by Maxwell’s equations. The electric and magnetic field to the material elements are determined by the 4-velocities of the elements. Considering the cavity as a vacuum, we impose the condition of vanishing spatial velocity and EM current. In cases where the material acts as a perfect conductor, the electric field is constrained to be orthogonal to the surface at the boundary.
@@
For a concise discussion, we utilize covariant perturbation theory. Within this framework, we introduce the lemma that a linear perturbation is gauge-invariant if and only if the Lie derivative of its leading-order value vanishes along all directions. The perturbation of Einstein’s equations transforms into the wave equation for the Riemann tensor, where its wave solutions manifest as GWs. Utilizing these solutions, we present the perturbation of the timelike geodesic congruence. The perturbations of spatial velocity and strain for the material exhibit homogeneous and inhomogeneous wave equations, respectively, as dictated by their equations of motion. Perturbing Maxwell’s equations yield the wave equation for the perturbation of the electric field. Within the cavity, this equation indicates that GWs contribute to the perturbation of the electric field solely through the coupling between the EM field and the Riemann tensor. Lastly, we point out that fixing the 4-velocity of the material in a coordinate system or considering a rigid body within context of general relativity is not appropriate.
@@
We solve Maxwell’s equations within the cavity. Using the solution for the electric field, we define the transfer function and the dimensionless GW signal. To derive the antenna pattern, we separate the GW signal into its strength and pattern function. Averaging the pattern function over the polarization angle yields the antenna pattern. As an illustrative example, we present the antenna pattern for the EM cavity using the TM010 mode at the resonance frequency.
@@
Our work enables a profound understanding of the operational principles of GW detectors employing EM cavities by clearly elucidating the interplay among electric fields, acoustic oscillations, and GWs. Due to the generality of our analysis, it is relevant to various GW detectors employing EM fields. Our paper does not address the measurement principles and methods for the induced EM fields. We perceive this as a challenging problem, and it will be the focus of our future work.
@@ Acknowledgements.
The authors thank Jai-chan Hwang and Sung Mook Lee for their helpful discussion. We appreciate APCTP for its hospitality during the completion of this work. This work was supported by IBS under the project code, IBS-R018-D1 and IBS-R017-D1-2023-a00. Y.-B.B. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. NRF-2021R1F1A1051269).

### v2

@@ V Summary and discussion
We introduce physical laws in a globally hyperbolic spacetime described by Einstein’s equations. The contracted Bianchi identities imply the conservation of stress–energy tensor. The congruence of the timelike geodesics without vorticity, have their spatial distribution determined by the extrinsic curvature. However, material elements, despite being influenced by gravity, deviate from geodesics due to elasticity stemming from their deformation. To account for the acoustic oscillation of the material, we introduce the four-velocity and strain, both contributing the stress–energy tensor. The equations of motion for these material elements encompass the temporal symmetry of the material metric, energy conservation, and the conservation of stress–energy tensor. The electromagnetic field is governed by Maxwell’s equations. The electric field and magnetic field to the material elements are determined by the four-velocities of the elements. Considering the cavity as a vacuum, we impose the condition of vanishing spatial velocity and EM current. In cases where the material acts as a perfect conductor, the electric field is constrained to be orthogonal to the surface at the boundary.
@@
For a concise discussion, we utilize covariant perturbation theory. Within this framework, we introduce the lemma that a linear perturbation is gauge invariant if and only if the Lie derivative of its leading-order value vanishes along all directions. The perturbation of Einstein’s equations transforms into the wave equation for the Riemann tensor, where its wave solutions manifest as GWs. Utilizing these solutions, we present the perturbation of the timelike geodesic congruence. The perturbations of spatial velocity and strain for the material exhibit homogeneous and inhomogeneous wave equations, respectively, as dictated by their equations of motion. Perturbing Maxwell’s equations yield the wave equation for the perturbation of the electric field. Within the cavity, this equation indicates that GWs contribute to the perturbation of the electric field solely through the coupling between the EM field and the Riemann tensor. Lastly, we point out that fixing the four-velocity of the material in a coordinate system or considering a rigid body within the context of general relativity is not appropriate.
@@
We solve Maxwell’s equations within the cavity. Using the solution for the electric field, we define the transfer function and the dimensionless GW signal. To derive the antenna pattern, we separate the GW signal into its strength and pattern function. Averaging the pattern function over the polarization angle yields the antenna pattern. As illustrative examples, we present the antenna patterns for the EM cavities using the $\mathrm{TM}_{010}$ and $\mathrm{TE}_{212-}$ modes, respectively, at their own resonance frequencies.
@@
Our work enables a profound understanding of the operational principles of GW detectors employing EM cavities by clearly elucidating the interplay among electric fields, acoustic oscillations, and GWs. Owing to the generality of our analysis, it is relevant to various GW detectors employing EM fields. Our paper does not address the measurement principles and methods for the induced EM fields. We perceive this as a challenging problem, and it will be the focus of our future work.
@@ Acknowledgements.
The authors thank Jai-chan Hwang and Sung Mook Lee for their helpful discussion. We appreciate APCTP for its hospitality during the completion of this work. This work was supported by IBS under the Project Codes, No. IBS-R018-D1 and No. IBS-R017-D1-2023-a00. Y.-B.B. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. NRF-2021R1F1A1051269). C.P. was supported by the NRF funded by the Korean government (No. NRF-2021M3F7A1082056 and No. NRF-2021R1A2C2012473). All authors contributed equally to this work.